# AIME: AI System Optimization via Multiple LLM Evaluators

### v1

@@ 1 Introduction
While prior art has compared the abilities of a single LLM for evaluations against those of multiple LLMs (Kocmi & Federmann, 2023; Ankner et al., 2024), in AI system optimization literature, there has been a lack of studies questioning the capabilities of using a single LLM evaluator to drive the optimization process. Recently, Yuksekgonul et al. (2024) has viewed the evaluation as a text-based analogy to the objective function for backpropagation (Hinton et al., 2006; Rumelhart et al., 1986) in deep learning optimization. The objective function is a crucial element in optimizing machine learning models (Christiano et al., 2017; Mescheder et al., 2018; Chakraborty et al., 2023; Kingma & Welling, 2014). This importance motivates us to analyze and strengthen the evaluation protocol of state-of-the-art (SoTA) AI system optimization frameworks by addressing a critical research question: What are the failure cases or tasks of utilizing only one LLM-based evaluator for text-based AI system optimization?.
@@
Figure 1: AI System Optimization Pipeline and Increased Error Detection and Success Rate with AIME-based Evaluation: [LEFT] Text-based AI system optimization with SoTA framework (Yuksekgonul et al., 2024) using our multiple LLM evaluator approach AIME (orange) and with single-evaluator approach (blue). [TOP RIGHT] The single-evaluator approach cannot detect an error in the generated code that fails all test cases. However, one of the evaluators of AIME could because the logical evaluator was independent from the correctness evaluator. [BOTTOM RIGHT] AIME-based optimization achieves a $\sim 16\%$ higher success rate than a single-evaluator approach in code generation tasks.
@@
  • Empirical Performance Over Single-Evaluation Approach: Using popular code generation dataset, LeetCodeHard (Shinn et al., 2024) and HumanEval (Chen et al., 2021), we perform an extensive study showing the superior prowess of AIME with $6$ evaluators over single evaluation to detect errors, with AIME achieve up to $62\%$ higher error detection rate than single evaluation. We then show that AIME-based optimization achieves up to a $16\%$ higher success rate on test cases than optimization with only a single evaluator. We also reveal that the choice of the number of evaluators and the combination of criteria to utilize can affect the success rate by up to $12\%$, emphasizing the design of AIME-based optimization is non-trivial.
@@ 2 Text-based AI System Optimization
where we first expand upon the sub-optimality in evaluation and then upper-bound using the total variation distance (sriperumbudur2009integral). We see that the term $d_{\text{TV}}(\pi_{e}^{*}(\cdot|x,y),\pi(\cdot|x,y))$ is fixed and it can’t be improved once we have the evaluator $\pi$. This result shows the hardness of a single evaluator reaching $\pi_{e}^{*}$ due to this constant gap and it will only reduce if our current LLM evaluator is near-optimal which is not true in majority of the scenarios. Empirically, Figure 1 demonstrates a practical observation where a single evaluator lets code errors go undetected, causing a large suboptimality gap from oracle performance in code generation tasks.
@@ 4 Experiments and Results
AIME and Single-Eval Implementation Details: We use TextGrad from Yuksekgonul et al. (2024) to implement AIME and Single-Eval. We chose TextGrad because it separates the evaluation and feedback into two separate LLM calls, making it better to analyze the evaluation module in isolation. In TextGrad, the system prompt that generates the initial code, $p_{\text{init}}$, is different from the system prompt that updates the code in the following refinement iterations $p_{\text{update}}$. At $t=0$, $p_{\text{init}}$ specifies to the LLM that it is a code generator while the $p_{\text{update}}$ from $1\leq t\leq T$ specifies that it generates a new version $y_{t+1}$ given the current code $y_{t}$ and the feedback $f_{t}$. The transition from $p_{\text{init}}$ to $p_{\text{update}}$ is explicitly programmed and not caused by the optimization process. Because the scope of this paper lies within the evaluation protocol, our AI system is a single LLM generator.
@@ 5 Related Works
AI System Optimization: Many prior works have studied the optimization of complex AI systems. madaan2024self was one of the first works to propose a text-based iterative feedback loop for refining LLMs, and Pryzant et al. (2023) established text-based gradients, or Textual Gradients, as feedback to an AI system. DSPy (Khattab et al., 2024; 2022; Singhvi et al., 2023), Trace Cheng et al. (2024), and TextGrad (Yuksekgonul et al., 2024) have formulated LLM and AI-based systems as a network of multiple layers and provided methods to optimize these system analogous to backpropagation and autodifferentiation. Chakraborty et al. (2024a); Ding et al. (2024) used a bi-level optimization formulation to align AI agents and systems. Text-based reinforcement learning has also been used to improve LLM-based systems (Shinn et al., 2024). Decoding and RLHF is an alternative method to optimize or align an LLM with gradient descent (Chakraborty et al., 2024b; Mudgal et al., 2023; Chakraborty et al., 2024c). While these works have shown tremendous results, there has been a gap in the literature we aim to address analyzing the effect of using multiple independent evaluations to optimize the AI system for a complex task, code generation (Chen et al., 2024; Zeng et al., 2024; Zhang et al., 2023; Jha et al., 2010; Shinn et al., 2024; Yuksekgonul et al., 2024; Zan et al., 2022; Jiang et al., 2024; Chen et al., 2021; Gulwani, 2010).

### v2

@@ Abstract
Bhrij Patel1,   Souradip Chakraborty1,   Wesley A. Suttle2,   Mengdi Wang3,
Amrit Singh Bedi*4,   Dinesh Manocha*1
1 University of Maryland, College Park
2 U.S. Army Research Laboratory, MD, USA
3 Princeton University
4 University of Central Florida
*Denotes Equal Advising
@@ 1 Introduction
While prior art has compared the abilities of a single LLM for evaluations against those of multiple LLMs (Kocmi & Federmann, 2023; Ankner et al., 2024), in AI system optimization literature, there has been a lack of studies questioning the capabilities of using a single LLM evaluator to drive the optimization process. Recently, Yuksekgonul et al. (2024) has viewed the evaluation as a text-based analogy to the objective function for backpropagation (Hinton et al., 2006; Rumelhart et al., 1986) in deep learning optimization. The objective function is a crucial element in optimizing machine learning models (Christiano et al., 2017; Mescheder et al., 2018; Chakraborty et al., 2023; Kingma & Welling, 2014). This importance motivates us to analyze and strengthen the evaluation protocol of state-of-the-art (SoTA) AI system optimization frameworks by addressing a critical research question: What are the failure cases or tasks of utilizing only one LLM-based evaluator for text-based AI system optimization?
@@
Figure 1: AI System Optimization Pipeline and Increased Error Detection and Success Rate with AIME-based Evaluation: [LEFT] Text-based AI system optimization with SoTA framework (Yuksekgonul et al., 2024) using our multiple LLM evaluator approach AIME (orange) and with single-evaluator approach (blue). [TOP RIGHT] The single-evaluator approach cannot detect an error in the generated code that fails all test cases. However, one of the evaluators of AIME could because the logical evaluator was independent from the correctness evaluator. [BOTTOM RIGHT] AIME-based optimization achieves $\sim 16\%$ higher success rate than a single-evaluator approach in code generation tasks.
@@
  • Empirical Performance Over Single-Evaluation Approach: Using popular code generation dataset, LeetCodeHard (Shinn et al., 2024) and HumanEval (Chen et al., 2021), we perform an extensive study showing the superior prowess of AIME with $6$ evaluators over single evaluation to detect errors, with AIME achieve up to $62\%$ higher error detection rate than single evaluation. We then show that AIME-based optimization achieves up to a $16\%$ higher success rate on test cases than optimization with only a single evaluator. We also reveal that the choice of the number of evaluators and the combination of criteria to utilize can affect the success rate by up to $12\%$, emphasizing the design of AIME-based optimization is non-trivial. We provide an anonymized code repository.  [^1] Repository to code: https://github.com/Bridge00/aime
@@ 2 Text-based AI System Optimization
where we first expand upon the sub-optimality in evaluation and then upper-bound using the total variation distance (Sriperumbudur et al., 2009). We see that the term $d_{\text{TV}}(\pi_{e}^{*}(\cdot|x,y),\pi(\cdot|x,y))$ is fixed and it cannot be improved once we have the evaluator $\pi$. This result shows the hardness of a single evaluator reaching $\pi_{e}^{*}$ due to this constant gap and it will only reduce if our current LLM evaluator is near-optimal which is not true in majority of the scenarios. Empirically, Figure 1 demonstrates a practical observation where a single evaluator lets code errors go undetected, causing a large suboptimality gap from oracle performance in code generation tasks.
@@ 4 Experiments and Results
AIME and Single-Eval Implementation Details: We use TextGrad from Yuksekgonul et al. (2024) to implement AIME and Single-Eval. We chose TextGrad because it separates the evaluation and feedback into two separate LLM calls, making it better to analyze the evaluation module in isolation. In TextGrad, the system prompt that generates the initial code, $p_{\text{init}}$, is different from the system prompt that updates the code in the following refinement iterations $p_{\text{update}}$. At $t=0$, $p_{\text{init}}$ specifies to the LLM that it is a code generator while the $p_{\text{update}}$ from $1\leq t\leq T$ specifies that it generates a new version $y_{t+1}$ given the current code $y_{t}$ and the feedback $f_{t}$. The transition from $p_{\text{init}}$ to $p_{\text{update}}$ is explicitly programmed and not caused by the optimization process. Because the scope of this paper lies within the evaluation protocol, our AI system is a single LLM generator.  [^2] We repeat the link to the repository: https://github.com/Bridge00/aime
@@ 5 Related Works
AI System Optimization: Many prior works have studied the optimization of complex AI systems. Madaan et al. (2024) was one of the first works to propose a text-based iterative feedback loop for refining LLMs, and Pryzant et al. (2023) established text-based gradients, or Textual Gradients, as feedback to an AI system. DSPy (Khattab et al., 2024; 2022; Singhvi et al., 2023), Trace Cheng et al. (2024), and TextGrad (Yuksekgonul et al., 2024) have formulated LLM and AI-based systems as a network of multiple layers and provided methods to optimize these system analogous to backpropagation and autodifferentiation. Chakraborty et al. (2024a); Ding et al. (2024) used a bi-level optimization formulation to align AI agents and systems. Text-based reinforcement learning has also been used to improve LLM-based systems (Shinn et al., 2024). Decoding and RLHF is an alternative method to optimize or align an LLM with gradient descent (Chakraborty et al., 2024b; Mudgal et al., 2023; Chakraborty et al., 2024c). While these works have shown tremendous results, there has been a gap in the literature we aim to address analyzing the effect of using multiple independent evaluations to optimize the AI system for a complex task, code generation (Chen et al., 2024; Zeng et al., 2024; Zhang et al., 2023; Jha et al., 2010; Shinn et al., 2024; Yuksekgonul et al., 2024; Zan et al., 2022; Jiang et al., 2024; Chen et al., 2021; Gulwani, 2010).